OpenStudy (anonymous):

PLEASE TEACH ME HOW TO DO IT!! What is 1.2% of 50? 2.4 4.2 0.6 6

hartnn (hartnn):

simple way is x% = x/100 so, 1.2% of 5 will be 1.2/100 * 50 =... ?

hartnn (hartnn):

1.2% of 50 will be 1.2/100 * 50 =... ?

OpenStudy (anonymous):

THANKS!

OpenStudy (amistre64):

the word "percent" mean "out of 100". the simplest way to do this is to divide 50 by 100, that gives you 1% then mutliply that by how many %s you need, in this case 1.2

hartnn (hartnn):

welcome :) if you want to verify the answer, tell us what u got.

OpenStudy (anonymous):

0.6?

OpenStudy (anonymous):

DECIMAL...

hartnn (hartnn):

yes, 0.6 is correct :)